An organic compound is found to contain 89.94% C and 10.06% H. It has a Molar Mass of 320g/mole. What is the empirical and true formula?


1
Expert's answer
2021-03-01T02:39:44-0500

w=n*Ar/Mr

n=w*Mr/Ar

n(C)=0.8994*320/12=24

n(H)=0.1006*320/1=32.

So, the true formula is C24H32, and the empirical formula is C3H4.